OpenStudy (anonymous):

what is the inverse of f(x)=x^3+1

OpenStudy (anonymous):

y = x^3 + 1. inverse y and x. x = y^3 + 1 ==> y^3 = x-1 ==> y = cube root of (x-1)

OpenStudy (anonymous):

cube root of (x-1)

OpenStudy (anonymous):

my book says the anser is \[5\sqrt{x-1}\]

OpenStudy (anonymous):

5th root? that is incorrect, unless f(x) is x^5

OpenStudy (anonymous):

no, i think the 5 is outside the square root

OpenStudy (anonymous):

Nope